# Finite temperature detection of quantum critical points via internal quantum teleportation

### v1

@@ I Introduction
Many important quantum information theory quantities were shown to be excellent quantum critical point (QCP) detectors at $T=0$ wu04 ; oli06 ; dil08 ; sar08 ; fan14 ; gir14 ; hot08 ; hot09 ; ike23a ; ike23b ; ike23c ; ike23d . However, it is known that some of them do not work properly when $T>0$. For example, the entanglement of formation woo98 between a pair of spins that belongs to a spin chain is zero before, at, and after the QCP if the chain is prepared above a certain threshold temperature wer10 ; wer10b . Therefore, both from a theoretical point of view and from practical aspects (we cannot cool a many-body system to $T=0$ due to the third law of thermodynamics), it is important to develop robust tools to characterize QPTs at finite $T$. In this case thermal fluctuations are present and we must have tools to detect QCPs that properly work in this scenario.
@@
In Ref. pav23 we went one step further and presented QCP detection tools that keep the most useful features of QD in detecting QCPs at finite $T$ but do not have its theoretical and experimental handicaps outlined above. These tools are based on the quantum teleportation protocol ben93 ; yeo02 ; rig17 and are scalable to high dimensional systems, have a clear experimental meaning, and work at temperatures where other tools, such as the entanglement of formation, already fail to spotlight QCPs.
In this work we show, under very general conditions, that we can simplify the tools to detect QCPs at finite $T$ given in Ref. pav23 , while still keeping their efficiency, robustness, and scalability. In the present proposal, we do not need an external qubit from the system to implement the teleportation based QCP detector. We even do not need to repeat the procedure using several different external qubits covering the whole Bloch sphere in order to detect a QCP pav23 . In the present proposal, the input state to be teleported is now fixed and belongs to the many-body system itself (see Fig. 1).
@@ II.1 General settings
The second measure we will be using in this work to quantify the similarity between two mixed states is the trace distance nie00 ,
@@ II.2 Specific settings
with $Z=Tr[e^{-H/kT}]$ being the partition function and $k$ the Boltzmann’s constant. The density matrix describing a pair of nearest neighbor spins, for instance $\rho_{23}$, is obtained tracing out from $\varrho$ all but spins $2$ and $3$. For all the models investigated here we have wer10b ; pav23 ; osb02
@@
where $\alpha=x,y,z$, can be found in Refs. yan66 ; clo66 ; klu92 ; bor05 ; boo08 ; tri10 ; tak99 ; lie61 ; bar70 ; bar71 ; pfe70 ; zho10 . In Ref. wer10b these calculations are reviewed and written in the present notation. The behavior of $\left\langle\sigma_{j}^{z}\right\rangle$ and $\left\langle\sigma_{j}^{\alpha}\sigma_{j+1}^{\alpha}\right\rangle$ for several values of $T$ as we change the tuning parameter for each one of the models given in Secs. III and IV are shown in Ref. pav23 .
@@
$\rho_{1}=\left(\begin{array}[]{cc}\frac{1+\left\langle\sigma^{z}\right\rangle}{2}&0\\ 0&\frac{1-\left\langle\sigma^{z}\right\rangle}{2}\end{array}\right),$ (28)
@@
where we have dropped the subscript from the one-point correlation function for simplicity.
Inserting Eqs. (1), (3), (4), (20), and (28) into Eq. (7) gives
@@ III The XXZ model in an external field
Looking at Fig. 3 we note that the two QCPs ($\Delta_{1}$ and $\Delta_{2}$) are clearly detected by discontinuities in the derivatives of $\overline{\mathcal{D}}$ with respect to $\Delta$ as we go through the critical points. Also, in contrast to the behavior of $\overline{\mathcal{F}}$ in the external teleportation approach pav23 , we do not see two extra discontinuities in the derivatives of $\overline{\mathcal{F}}$ and of $\overline{\mathcal{D}}$ [cf. Figs. 2 and 3 with the corresponding ones given in Ref. pav23 ]. In the present case, we only have one small kink between the two QCPs.
@@ IV.1 The $\lambda$ transition
In Figs. 6, 7, and 8 we plot the maximum mean fidelity $\overline{\mathcal{F}}$ as a function of $\lambda$ for three values of $\gamma$ and for several values of temperature. The chosen values for $\gamma$ are such that we obtain the isotropic XX model ($\gamma=0.0$), the anisotropic XY model ($\gamma=0.5$), and the Ising model ($\gamma=1.0$).
@@
When $\gamma=0.5$ (anisotropic XY model) and $\gamma=1.0$ (Ising transverse model), the QCP is determined, respectively, by the inflection points of $\overline{\mathcal{F}}$ and $d\overline{\mathcal{F}}/d\lambda$ that occur exactly at $\lambda_{c}=1.0$ when $T=0$ (see Figs. 7 and 8). For higher values of $T$, the inflection points are displaced away from $\lambda_{c}$ and the best strategy to estimate the QCP is by determining the location of the maximum (minimum) of the second order derivatives of $\overline{\mathcal{F}}$ with respect to $\lambda$. The location of these extremum values lie close to $\lambda_{c}$ when $kT\lesssim 0.1$ and, as we will show below, by extrapolating to the absolute zero temperature we can predict its correct value. Note also that the one- and two-point correlation functions have inflection points at the QCP when $T=0$ pav23 .
In Figs. 9, 10, and 11 we plot the minimum mean trace distance $\overline{\mathcal{D}}$ as a function of $\lambda$ for the same three values of $\gamma$ and temperatures shown in Figs. 6 to 8. Looking at Figs. 6 to 11, and taking into account that a high value for $\overline{\mathcal{F}}$ implies a low value for $\overline{\mathcal{D}}$, we realize that the behavior of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ about the QCPs are very similar for a given value of $\gamma$.
@@
The same qualitative features observed for the behavior of $\overline{\mathcal{F}}$ as a function of $\lambda$ can be seen for $\overline{\mathcal{D}}$ when we work with $\gamma=0.5$ (anisotropic XY model) and $\gamma=1.0$ (Ising transverse model). Again, at $T=0$ the critical point $\lambda_{c}$ is given, respectively, by an inflection point of $\overline{\mathcal{D}}$ and of $d\overline{\mathcal{D}}/d\lambda$ at $\lambda_{c}=1.0$ (see Figs. 10 and 11). If we increase $T$, though, the inflection points move away from $\lambda_{c}=1.0$. As we will see, using finite $T$ data we can best estimate the correct value for $\lambda_{c}$ by computing the location of the extremum values of the second order derivatives of $\overline{\mathcal{D}}$ with respect to $\lambda$ and then numerically taking the $T=0$ limit.
@@
We can be more quantitative in the prediction of the location of the QCP by following the same strategy of Ref. pav23 that was already explained in Sec. III. For all the quantities shown in Figs. 12 to 14, we excluded the $kT=0.0$ point and implemented a linear regression with the remaining data ($kT=0.01,0.02,\ldots,0.10$). Using the fitted curves, we took the $kT=0$ limit and obtained for all cases and within the numerical errors already reported the correct value for the QCP.
@@
Figure 15: (color online) Mean fidelities $\overline{F}(S_{\Psi^{\pm}})$ and $\overline{F}(S_{\Phi^{\pm}})$ [Eqs. (31) and (32)] and mean trace distances $\overline{D}(S_{\Psi^{\pm}})$ and $\overline{D}(S_{\Phi^{\pm}})$ [Eqs. (36) and (37)] as a function of $\lambda$ when $kT=0$ and $\gamma=0.0$ (XX model in a transverse field). The highlighted points, which we called “crossing points”, are located exactly at the cusps of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ seen in Figs. 6 and 9, respectively. Note that these cusps are not related to QPTs and they move to higher values of $\lambda$ as we increase the temperature.
Plotting the same curves for $\gamma=0.5$ and $\gamma=1.0$, we realize that the mean fidelity curves as well as the mean trace distance ones do not cross each other. This is why we do not have a cusp in Figs. 7 and 8 and also in Fig. 11. We do have a small cusp, though, for the curves of $\overline{\mathcal{D}}$ given in Fig. 10, where we fixed $\gamma=0.5$. This cusp is not related to a QPT and we can understand its origin by carefully looking at the functional form of $\overline{D}(S_{\Phi^{\pm}})$, which gives the minimum mean trace distance $\overline{\mathcal{D}}$ when $0\leq\lambda\leq 2$ (see Fig. 16).

### v2

@@ I Introduction
Many important quantum information theory quantities were shown to be excellent quantum critical point (QCP) detectors at $T=0$ ost02 ; osb02 ; wu04 ; oli06 ; dil08 ; sar08 ; fan14 ; gir14 ; hot08 ; hot09 ; ike23a ; ike23b ; ike23c ; ike23d . However, it is known that some of them do not work properly when $T>0$. For example, the entanglement of formation woo98 between a pair of spins that belongs to a spin chain is zero before, at, and after the QCP if the chain is prepared above a certain threshold temperature wer10 ; wer10b . Therefore, both from a theoretical point of view and from practical aspects (we cannot cool a many-body system to $T=0$ due to the third law of thermodynamics), it is important to develop robust tools to characterize QPTs at finite $T$. In this case thermal fluctuations are present and we must have tools to detect QCPs that properly work in this scenario.
@@
In Ref. pav23 we went one step further and presented QCP detection tools that keep the most useful features of QD in detecting QCPs at finite $T$ but do not have its theoretical and experimental handicaps outlined above. These tools are based on the quantum teleportation protocol ben93 ; yeo02 ; rig17 and are scalable to high dimensional systems, have a clear experimental meaning, and work at temperatures where other tools, such as the entanglement of formation, already fail to spotlight QCPs. See also Refs. ven06 ; ven07 ; ven07b ; aba23 for other applications of the quantum teleportation protocol in spin chains, in particular the teleportation of a quantum state between distant spins of the chain.
In this work we show, under very general conditions, that we can simplify the tools to detect QCPs at finite $T$ given in Ref. pav23 , while still keeping their efficiency, robustness, and scalability. In the present proposal, we do not need an external qubit from the system to implement the teleportation based QCP detector. We even do not need to repeat the procedure using several different external qubits covering the whole Bloch sphere in order to detect a QCP pav23 . In the present proposal, the input state to be teleported is now fixed and belongs to the many-body system itself (see Fig. 1). Also, and similarly to the tool developed in Ref. pav23 , we do not need to know the order parameter associated with the QPT to implement it (see Refs. wer10b ; pav23 for more details).
@@ II.1 General settings
The second measure we will be using in this work to quantify the similarity between two mixed states is the trace distance nie00 ; tos23 ; tos24 ,
@@ II.2 Specific settings
with $Z=Tr[e^{-H/kT}]$ being the partition function and $k$ the Boltzmann’s constant.
The density matrix describing three nearest neighbors from the spin chain is given by $\rho_{123}$, where it is obtained tracing out from $\varrho$ all but spins $1,2$ and $3$. Note that the initial state $\rho_{123}$ can be obtained using any three nearest neighbors from the spin chain. After putting the whole chain in contact with a heat reservoir at temperature $T$ and waiting for the system to reach the thermodynamic equilibrium, any three consecutive spins of the chain can be used to implement the internal teleportation protocol as described below.
The density matrix $\rho_{123}$ is in general not given by Eq. (1), where $\rho_{1}$ is given by tracing out from $\varrho$ all but spin $1$ and $\rho_{23}$ is the state obtained by tracing out from $\varrho$ all but spins $2$ and $3$. The state $\rho_{1}$ describes a single spin from the chain and $\rho_{23}$ any nearest neighbor pair of spins from the chain. We have, therefore, to prepare an ensemble from $\rho_{123}$ that is effectively given by Eq. (1). Before we explain how this can be done, we first give the explicit forms of $\rho_{23}$ and $\rho_{1}$ in the standard basis.
For all the models investigated here we have wer10b ; pav23 ; osb02
@@
where $\alpha=x,y,z$, can be found in Refs. yan66 ; clo66 ; klu92 ; bor05 ; boo06 ; boo08 ; tri10 ; tak99 ; lie61 ; bar70 ; bar71 ; pfe70 ; zho10 . In Ref. wer10b these calculations are reviewed and written in the present notation. The behavior of $\left\langle\sigma_{j}^{z}\right\rangle$ and $\left\langle\sigma_{j}^{\alpha}\sigma_{j+1}^{\alpha}\right\rangle$ for several values of $T$ as we change the tuning parameter for each one of the models given in Secs. III and IV are shown in Ref. pav23 .
@@
$\rho_{1}=\left(\begin{array}[]{cc}p_{0}&0\\ 0&p_{1}\end{array}\right)=p_{0}|0\rangle\langle 0|+p_{1}|1\rangle\langle 1|,$ (28)
where
$p_{0}$ $=$ $\frac{1+\left\langle\sigma^{z}\right\rangle}{2},$ (29)
$p_{1}$ $=$ $\frac{1-\left\langle\sigma^{z}\right\rangle}{2}.$ (30)
@@
In Eqs. (29) and (30) we have dropped the subscript from the one-point correlation function for simplicity.
Using Eq. (28) and that
$\rho_{23}=Tr_{1}[\rho_{123}]=\,_{1}\langle 0|\rho_{123}|0\rangle_{1}+\,_{1}\langle 1|\rho_{123}|1\rangle_{1},$
the state (1) can be written as follows,
$\rho$ $=$ $p_{0}|0\rangle_{1}\,{}_{1}\langle 0|\otimes\,_{1}\langle 0|\rho_{123}|0\rangle_{1}$ (31)
$+p_{0}|0\rangle_{1}\,{}_{1}\langle 0|\otimes\,_{1}\langle 1|\rho_{123}|1\rangle_{1}$
$+p_{1}|1\rangle_{1}\,{}_{1}\langle 1|\otimes\,_{1}\langle 0|\rho_{123}|0\rangle_{1}$
$+p_{1}|1\rangle_{1}\,{}_{1}\langle 1|\otimes\,_{1}\langle 1|\rho_{123}|1\rangle_{1}.$
Writing $\rho$ as given by Eq. (31) readily suggests the steps Alice needs to do to effectively obtain state (1). She projects qubit 1 in the computational basis. She either obtains $|0\rangle_{1}$ or $|1\rangle_{1}$ with probabilities $p_{0}$ and $p_{1}$, respectively. According to the measurement postulate of quantum mechanics nie00 , the state $\rho_{123}$ changes in each case to
$\rho_{123}\rightarrow\frac{P_{|0\rangle}\rho_{123}P_{|0\rangle}}{p_{0}}=|0\rangle_{1}\,{}_{1}\langle 0|\otimes\frac{\,{}_{1}\langle 0|\rho_{123}|0\rangle_{1}}{p_{0}},$ (32)
$\rho_{123}\rightarrow\frac{P_{|1\rangle}\rho_{123}P_{|1\rangle}}{p_{1}}=|1\rangle_{1}\,{}_{1}\langle 1|\otimes\frac{\,{}_{1}\langle 1|\rho_{123}|1\rangle_{1}}{p_{1}},$ (33)
where $P_{|j\rangle}=|j\rangle_{1}\,{}_{1}\langle j|$, with $j=0,1$, is the projector onto the state $|j\rangle_{1}$.
Also, if after measuring her qubit Alice applies the spin flip unitary operation, namely, $\sigma_{1}^{x}|0(1)\rangle=|1(0)\rangle$, we have for each possible measurement result,
$\rho_{123}\rightarrow\frac{\sigma_{1}^{x}P_{|0\rangle}\rho_{123}P_{|0\rangle}\sigma_{1}^{x}}{p_{0}}=|1\rangle_{1}\,{}_{1}\langle 1|\otimes\frac{\,{}_{1}\langle 0|\rho_{123}|0\rangle_{1}}{p_{0}},$ (34)
$\rho_{123}\rightarrow\frac{\sigma_{1}^{x}P_{|1\rangle}\rho_{123}P_{|1\rangle}\sigma_{1}^{x}}{p_{1}}=|0\rangle_{1}\,{}_{1}\langle 0|\otimes\frac{\,{}_{1}\langle 1|\rho_{123}|1\rangle_{1}}{p_{1}}.$ (35)
From four subensembles, each one with an equal number $N$ of systems described by Eqs. (32), (33), (34), or (35), Alice builds the following ensemble. She picks $(p_{0})^{2}N$ states given by (32), $(p_{1})^{2}N$ states given by (33), $p_{0}p_{1}N$ states given by (34), and $p_{0}p_{1}N$ states given by (35). Combining these four subensembles leads to $N$ states described by Eq. (31), or equivalently, Eq. (1).111 It is worth noticing that if Alice picks qubit $1$ sufficiently away from qubits $2$ and $3$, the state $\rho_{123}$ is already given by Eq. (1). Also, if we prepare two identical spin chains and pick qubit $1$ from one chain and a pair of nearest neighbors from the other chain, we also have Eq. (1) describing those three qubits.
Note that due to the linearity of all the steps related to the implementation of the teleportation protocol, from the experimental point of view Alice’s job to effectively obtain Eq. (1) is very simple. She just needs to measure the qubit $1$ in the computational basis and then either do nothing or implement the bit flip operation before initiating the standard teleportation protocol as described in Sec. II.1. Being more specific, if she measures $|0\rangle$ she does nothing to her qubit $(p_{0})^{2}$ of the times or she flips it $p_{0}p_{1}$ of the times before starting the teleportation protocol. If she measures $|1\rangle$, she does nothing $(p_{1})^{2}$ of the times or she flips it $p_{0}p_{1}$ of the times before implementing the teleportation protocol. Also, she must work with an equal amount of cases where she measured $|0\rangle$ or $|1\rangle$ to mimic the state given by Eq. (31) following the prescription given above. This is easily accomplished discarding the exceeding cases.
The bottom line is that by proceeding as described above and due to the linearity of all the steps involved in the execution of the teleportation protocol, the effective initial state will be given by Eq. (1) and Bob’s qubit at the end of several runs of the teleportation protocol will effectively be described by Eq. (2) whenever Alice measures the Bell state $j$. Figure 2 illustrates all steps of a single run of the protocol.
Figure 2: (color online) One run of the internal teleportation protocol can be described as follows. First, Alice projects the qubit to be teleported (qubit $1$) onto the computational basis [panel (a)]. Second, she either does nothing with her qubit or applies onto it the spin flip operation ($\sigma_{1}^{x}$) according to the prescription given in the text [panel (b)]. Third, Alice and Bob execute the standard teleportation protocol following the steps given in Fig. 1 [panels (c) and (d)].
Alternatively one could use the state $\rho_{123}$ to theoretically simulate the execution of the internal teleportation protocol. The state $\rho_{123}$ is a $8\times 8$ matrix depending not only on one- and two-point correlation functions but also on the genuine three-point ones ($\langle\sigma^{\alpha}_{j}\sigma^{\beta}_{j+1}\sigma^{\gamma}_{j+1}\rangle$). Nevertheless, due to the factorization property of the correlation functions of the XXZ spin chain boo06 , the three-point correlations for this model can be expressed very non-trivially in terms of the one- and two-point correlations. However, the computation of the three-point correlation functions for finite temperature and for the other models here investigated is a quite involved task boo06 ; boo08 ; tri10 . Thus, in order to test the usefulness of the internal teleportation QCP detector, we worked with an initial state that depends at most on two-point correlation functions of two consecutive spin. It would be very interesting to extend the present analysis to the case where the state $\rho_{123}$ is used as the initial state. We believe that the internal teleportation QCP detector will also work in this case. Perhaps giving results even better than the ones reported here. We should also note that from the experimental point of view working with $\rho_{123}$ is even simpler. In this case Alice will neither need to project qubit $1$ onto the computational basis or flip her spin before implementing the standard teleportation protocol.
We now return to the computation of the relevant quantities needed to assess the efficiency of the teleportation protocol. Inserting Eqs. (1), (3), (4), (20), and (28) into Eq. (7) gives
@@ III The XXZ model in an external field
Looking at Fig. 4 we note that the two QCPs ($\Delta_{1}$ and $\Delta_{2}$) are clearly detected by discontinuities in the derivatives of $\overline{\mathcal{D}}$ with respect to $\Delta$ as we go through the critical points. Also, in contrast to the behavior of $\overline{\mathcal{F}}$ in the external teleportation approach pav23 , we do not see two extra discontinuities in the derivatives of $\overline{\mathcal{F}}$ and of $\overline{\mathcal{D}}$ [cf. Figs. 3 and 4 with the corresponding ones given in Ref. pav23 ]. In the present case, we only have one small kink between the two QCPs.
@@ IV.1 The $\lambda$ transition
In Figs. 7, 8, and 9 we plot the maximum mean fidelity $\overline{\mathcal{F}}$ as a function of $\lambda$ for three values of $\gamma$ and for several values of temperature. The chosen values for $\gamma$ are such that we obtain the isotropic XX model ($\gamma=0.0$), the anisotropic XY model ($\gamma=0.5$), and the Ising model ($\gamma=1.0$).
@@
When $\gamma=0.5$ (anisotropic XY model) and $\gamma=1.0$ (Ising transverse model), the QCP is determined, respectively, by the inflection points of $\overline{\mathcal{F}}$ and $d\overline{\mathcal{F}}/d\lambda$ that occur exactly at $\lambda_{c}=1.0$ when $T=0$ (see Figs. 8 and 9). For higher values of $T$, the inflection points are displaced away from $\lambda_{c}$ and the best strategy to estimate the QCP is by determining the location of the maximum (minimum) of the second order derivatives of $\overline{\mathcal{F}}$ with respect to $\lambda$. The location of these extremum values lie close to $\lambda_{c}$ when $kT\lesssim 0.1$ and, as we will show below, by extrapolating to the absolute zero temperature we can predict its correct value. Note also that the one- and two-point correlation functions have inflection points at the QCP when $T=0$ pav23 .
In Figs. 10, 11, and 12 we plot the minimum mean trace distance $\overline{\mathcal{D}}$ as a function of $\lambda$ for the same three values of $\gamma$ and temperatures shown in Figs. 7 to 9. Looking at Figs. 7 to 12, and taking into account that a high value for $\overline{\mathcal{F}}$ implies a low value for $\overline{\mathcal{D}}$, we realize that the behavior of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ about the QCPs are very similar for a given value of $\gamma$.
@@
The same qualitative features observed for the behavior of $\overline{\mathcal{F}}$ as a function of $\lambda$ can be seen for $\overline{\mathcal{D}}$ when we work with $\gamma=0.5$ (anisotropic XY model) and $\gamma=1.0$ (Ising transverse model). Again, at $T=0$ the critical point $\lambda_{c}$ is given, respectively, by an inflection point of $\overline{\mathcal{D}}$ and of $d\overline{\mathcal{D}}/d\lambda$ at $\lambda_{c}=1.0$ (see Figs. 11 and 12). If we increase $T$, though, the inflection points move away from $\lambda_{c}=1.0$. As we will see, using finite $T$ data we can best estimate the correct value for $\lambda_{c}$ by computing the location of the extremum values of the second order derivatives of $\overline{\mathcal{D}}$ with respect to $\lambda$ and then numerically taking the $T=0$ limit.
@@
We can be more quantitative in the prediction of the location of the QCP by following the same strategy of Ref. pav23 that was already explained in Sec. III. For all the quantities shown in Figs. 13 to 15, we excluded the $kT=0.0$ point and implemented a linear regression with the remaining data ($kT=0.01,0.02,\ldots,0.10$). Using the fitted curves, we took the $kT=0$ limit and obtained for all cases and within the numerical errors already reported the correct value for the QCP.
@@
Figure 16: (color online) Mean fidelities $\overline{F}(S_{\Psi^{\pm}})$ and $\overline{F}(S_{\Phi^{\pm}})$ [Eqs. (38) and (39)] and mean trace distances $\overline{D}(S_{\Psi^{\pm}})$ and $\overline{D}(S_{\Phi^{\pm}})$ [Eqs. (43) and (44)] as a function of $\lambda$ when $kT=0$ and $\gamma=0.0$ (XX model in a transverse field). The highlighted points, which we called “crossing points”, are located exactly at the cusps of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ seen in Figs. 7 and 10, respectively. Note that these cusps are not related to QPTs and they move to higher values of $\lambda$ as we increase the temperature.
Plotting the same curves for $\gamma=0.5$ and $\gamma=1.0$, we realize that the mean fidelity curves as well as the mean trace distance ones do not cross each other. This is why we do not have a cusp in Figs. 8 and 9 and also in Fig. 12. We do have a small cusp, though, for the curves of $\overline{\mathcal{D}}$ given in Fig. 11, where we fixed $\gamma=0.5$. This cusp is not related to a QPT and we can understand its origin by carefully looking at the functional form of $\overline{D}(S_{\Phi^{\pm}})$, which gives the minimum mean trace distance $\overline{\mathcal{D}}$ when $0\leq\lambda\leq 2$ (see Fig. 17).